A radio active substance of half life period 2 hrs is present in a reactor. At the initial moment the mass of the substance is 50 mg. Six hours later the mass of the radioactive substance would be nearly_________. 

Grade:10

1 Answers

govardhan
8 years ago
given that half life period=2 hours
INITIAL MASS=50 MG
GIVEN TIME =6HOURS
WE USE THIS RELATION N=N0/2^n
WHERE n=GIVEN TIME/HALF LIFE PERIOD
n=6/2
n=3
now
N=N0/2^n
N=50mg/2^3
N=50mg/8
N=6.25mg
Six hours later the mass of the radioactive substance would be nearly 6.25 mg.